# Evaluating Sparse Autoencoders: From Shallow Design to Matching Pursuit

### v1

@@ Abstract
Sparse autoencoders (SAEs) have recently become central tools for interpretability, leveraging dictionary learning principles to extract sparse, interpretable features from neural representations whose underlying structure is typically unknown. This paper evaluates SAEs in a controlled setting using MNIST, which reveals that current shallow architectures implicitly rely on a quasi-orthogonality assumption that limits the ability to extract correlated features. To move beyond this, we introduce a multi-iteration SAE by unrolling Matching Pursuit (MP-SAE), enabling the residual-guided extraction of correlated features that arise in hierarchical settings such as handwritten digit generation while guaranteeing monotonic improvement of the reconstruction as more atoms are selected.
@@ 1 Introduction
${\bm{x}}^{i}\approx{\bm{D}}{\bm{z}}^{i}\quad\text{subject to}\quad\|{\bm{z}}^{i}\|_{0}\leq k,\quad\forall\ i=1,\dots,n.\vspace{-1.5mm}$ (1)
@@
Sparsity has been established as a useful prior for interpretability (Mairal et al., 2014; Lipton, 2017; Ribeiro et al., 2016). Building on this principle, recent work has proposed the use of sparse autoencoders (SAEs) to extract human-interpretable features from the internal activations of large language models (LLMs) (Elhage et al., 2022; Cunningham et al., 2023; Bricken et al., 2023; Rajamanoharan et al., 2024; Gao et al., 2025). These methods are motivated by the linear representation hypothesis  (Arora et al., 2016; Olah et al., 2020; Park et al., 2024) and the superposition hypothesis  (Elhage et al., 2022), i.e., internal representations can be modeled as sparse linear combinations of semantic directions.
@@
Our Contributions In this work, we revisit the relationship between SAEs and dictionary learning by testing modern architectures on MNIST, a widely used benchmark in sparse coding (Aharon et al., 2006; Makhzani & Frey, 2014). Despite similarities in architecture, we find that SAEs with different sparsity mechanisms yield structurally distinct dictionaries. These differences may have important implications for interpretability.
@@ 2 Background
Orthogonality and Limitations of Shallow Recovery Sparse recovery (i.e., the inner problem optimization) in one iteration is theoretically guaranteed only when the dictionary ${\bm{D}}$ is sufficiently incoherent, i.e., when the mutual coherence $\mu({\bm{D}})=\max_{i\neq j}|{\bm{D}}_{i}^{\top}{\bm{D}}_{j}|$ is small (Makhzani & Frey, 2014; Arora et al., 2015). However, concepts underlying natural data may exhibit high coherence. This limits the ability of shallow, one-shot inference, sparse autoencoders to extract coherent concepts, and motivates the use of unrolled sparse-coding-based networks (Rambhatla et al., 2018; Tolooshams & Ba, 2022).
@@
3 Unrolling Matching Pursuit into MP-SAE
We propose Matching Pursuit Sparse Autoencoder (MP-SAE), an iterative inference procedure by unrolling Matching Pursuit (MP) (Mallat & Zhang, 1993) into a sparse autoencoder architecture. Unlike shallow SAEs, solving the inner problem in a single forward pass, MP-SAE performs inference sequentially.
Iterative Inference via Residual Updates  Matching Pursuit starts with a residual, defined as the difference between the input $\bm{x}$ and its current reconstruction $\hat{\bm{x}}$, i.e., $\bm{r}=\bm{x}-\hat{\bm{x}}$. Initially, this residual equals the input (or $\bm{x}-\bm{b}_{\text{pre}}$). MP-SAE iteratively reduces this residual by adding more atoms.
@@ 3 Unrolling Matching Pursuit into MP-SAE
Unlike TopK that chooses all activations at once, MP infer sequentially: it selects one atom at a time, removes its contribution from the residual, and continues. Because each selected atom is subtracted from the residual, the algorithm naturally explores new directions —each selection is driven by what remains unexplained, pushing the model toward atoms that are less redundant and more complementary. This residual-driven mechanism improves diversity in the selected features and enhances robustness in the presence of dictionary coherence. Indeed, MP can achieve exponential convergence and accurately recover active components even when the dictionary is highly coherent, as long as it satisfies a block-incoherence condition, where correlations are localized within small groups of atoms (Peotta & Vandergheynst, 2007).
@@ 4 Results
Learned Concepts In the following, we focus on dictionaries trained with $k=10$ and $p=1000$. The top row in Figure 3 shows 25 atoms with highest activation frequency for each SAE. All methods except ReLU appear to learn pen-stroke-like patterns, while ReLU learns atoms resembling full digits. The pen strokes learned by MP-SAE appear more precise. Interestingly, all shallow methods include a “negative” atom that closely resembles $-\bm{b}_{\text{pre}}$, which is the most frequently activated atom in ReLU and JumpReLU.
@@
This hierarchical behavior is supported by Figure 4(b), which shows that atoms with higher average activation values tend to be selected in the early layers of MP-SAE (the first iterations of Matching Pursuit). These atoms correspond to the digit-like patterns in the bottom row of Figure 3, and are refined by later atoms capturing more localized pen strokes (top row). This progression suggests a hierarchical structure in MP-SAE, building reconstructions from coarse to fine features (see Appendix 6).
Coherence Finally, we assess coherence for both the learned dictionary and the atoms selected at inference. To move beyond pairwise similarity captured by mutual coherence, the Babel function (Tropp, 2004) is used (see Equation 2). It measures cumulative coherence $\mu_{1}(r)$, offering a more comprehensive view of redundancy within the dictionary. $\mu_{1}(r)$ reflects how well a single atom can be approximated by a group of $r$ others; lower values indicate lower redundancy.
@@ 5 Conclusion
We introduce MP-SAE and show through small-scale experiments on MNIST that it improves expressivity, learns hierarchical features, and overcomes coherence limitations of shallow SAEs. These experiments reveal distinct representation behaviors across SAEs and offer a foundation for building more interpretable sparse autoencoders.

### v2

@@ Abstract
Sparse autoencoders (SAEs) have recently become central tools for interpretability, leveraging dictionary learning principles to extract sparse, interpretable features from neural representations whose underlying structure is typically unknown. This paper evaluates SAEs in a controlled setting using MNIST, which reveals that current shallow architectures implicitly rely on a quasi-orthogonality assumption that limits the ability to extract correlated features. To move beyond this, we compare them with an iterative SAE that unrolls Matching Pursuit (MP-SAE), enabling the residual-guided extraction of correlated features that arise in hierarchical settings such as handwritten digit generation while guaranteeing monotonic improvement of the reconstruction as more atoms are selected.
@@ 1 Introduction
${\bm{x}}^{i}\approx{\bm{D}}{\bm{z}}^{i}\quad\text{subject to}\quad\|{\bm{z}}^{i}\|_{0}\leq k,\quad\forall\ i=1,\dots,n.$ (1)
@@
Sparsity has been established as a useful prior for interpretability (Mairal et al., 2014; Lipton, 2017; Ribeiro et al., 2016). Building on this principle, recent work has proposed the use of sparse autoencoders (SAEs) to extract human-interpretable features from the internal activations of large language models (LLMs) (Elhage et al., 2022; Cunningham et al., 2023; Bricken et al., 2023; Rajamanoharan et al., 2024; Gao et al., 2025). These methods are motivated by the linear representation hypothesis  (Arora et al., 2016; Park et al., 2024; Elhage et al., 2022), which posits that internal representations can be modeled as sparse linear combinations of semantic directions. Unlike the classical dictionary learning methods described above, SAEs solve the inner problem in a single iteration.
@@
Our Contributions In this work, we revisit the relationship between SAEs and dictionary learning by testing common interpretability SAEs architectures on MNIST, a controlled and well-established dataset in sparse coding literature. Despite similarities in architecture, we find that SAEs with different sparsity mechanisms yield structurally distinct dictionaries. These differences may have important implications for interpretability.
@@ 2 Background
Orthogonality and Limitations of Shallow Recovery Sparse recovery (i.e., the inner problem optimization) in one iteration is theoretically guaranteed only when the dictionary $\bm{D}$} is sufficiently incoherent, i.e., when the mutual coherence $\mu({\bm{D}})=\max_{i\neq j}|{\bm{D}}_{i}^{\top}{\bm{D}}_{j}|$ is small (Makhzani & Frey, 2014; Arora et al., 2015). However, concepts underlying natural data may exhibit high coherence. This limitation motivates the use of unrolled sparse coding networks, which address the inner problem through multiple iterative updates, progressively refining the estimate of the sparse code $z$ as in classical sparse recovery methods. Shallow SAEs can be seen as approximating unrolled networks by performing only a single iteration: they rely on the same nonlinearities – JumpReLU implementing hard thresholding as in NOODL (Rambhatla et al., 2018), and ReLU implementing soft thresholding as in PUDLE (Tolooshams & Ba, 2022) – but do not perform iterative updates.
3 Unrolling Matching Pursuit into MP-SAE
We propose the Matching Pursuit Sparse Autoencoder (MP-SAE), which unrolls Matching Pursuit (MP) (Mallat & Zhang, 1993) into an iterative sparse autoencoder architecture. Previously used for image restoration tasks under the name Learned-MP (Khatib et al., 2021), our work focuses on the dictionary properties learned by MP-SAE and compares them to those obtained with shallow SAEs.
Iterative Inference via Residual Updates  MP-SAE can be seen as an unrolled variant of TopK. Instead of selecting $k$ atoms in a single iteration, as in TopK, it selects one atom at each of the $k$ iterations. This allows the model to subtract the contribution of each selected atom from the input before the next selection, so that subsequent atoms are chosen to explain what remains unexplained. This unexplained part is captured by the residual, defined as the difference between the input $\bm{x}$ and its current reconstruction $\hat{\bm{x}}$, i.e., $\bm{r}=\bm{x}-\hat{\bm{x}}$. Initially, the residual is equal to the input (or $\bm{x}-\bm{b}_{\text{pre}}$).
@@ 3 Unrolling Matching Pursuit into MP-SAE
Because each selected atom is subtracted from the residual at each iteration, the algorithm naturally explores new directions – each selection is driven by what remains unexplained, pushing the model toward atoms that are less redundant and more complementary. This residual-driven mechanism improves diversity in the selected features and enhances robustness in the presence of dictionary coherence. Indeed, MP can achieve exponential convergence and accurately recover active components even when the dictionary is highly coherent, as long as it satisfies a block-incoherence condition, where correlations are localized within small groups of atoms (Peotta & Vandergheynst, 2007).
@@ 4 Results
Learned Concepts In the following, we focus on dictionaries trained with $k=10$ and $p=1000$. The top row in Figure 3 shows 25 atoms with the highest activation frequency for each SAE. All methods except ReLU appear to learn pen-stroke-like patterns, while ReLU learns atoms resembling full digits. The pen strokes learned by MP-SAE appear more precise. Interestingly, all shallow methods include a “negative” atom that closely resembles $\bm{b}_{\text{pre}}$, which is the most frequently activated atom in ReLU and JumpReLU.
This hierarchical behavior is supported by Figure 4(b), which shows that atoms with higher average activation values tend to be selected in the early layers of MP-SAE (the first iterations of Matching Pursuit). These atoms correspond to digit-like patterns in the bottom row of Figure 3, and are refined by later atoms capturing more localized pen strokes (top row). This progression suggests a hierarchical structure in MP-SAE, reconstructing from coarse to fine features (see Appendix 6).
@@
Coherence Finally, we assess the coherence of both the learned dictionary and the atoms selected at inference time using the Babel score (Tropp, 2004). Unlike mutual coherence, which captures only the maximum pairwise similarity between atoms, the Babel function provides a more comprehensive measure of redundancy by quantifying cumulative coherence. Given a dictionary ${\bm{D}}=[{\bm{D}}_{1},\dots,{\bm{D}}_{p}]\in\mathbb{R}^{m\times p}$ with unit-norm columns, the Babel function of order $r$ is defined as:
$\mu_{1}(r)=\max_{\begin{subarray}{c}S\subset[p],|S|=r\end{subarray}}\left(\max_{j\notin S}\sum_{i\in S}\left|{\bm{D}}_{i}^{\top}{\bm{D}}_{j}\right|\right)$
The value $\mu_{1}(r)$ reflects how well a single atom can be approximated by a group of $r$ others—lower values indicate lower redundancy. Further details are provided in Appendix 7.
@@ 5 Conclusion
We introduce MP-SAE and show through small-scale experiments on MNIST that it improves expressivity, learns hierarchical features, and overcomes coherence limitations of shallow SAEs. These experiments also reveal distinct representation behaviors across SAEs and offer a foundation for building more interpretable sparse autoencoders.